### v2

# GLOBE: Accurate and Generalizable PDE Surrogates using Domain-Inspired Architectures and Equivariances

Abstract
We introduce GLOBE, a new neural surrogate for homogeneous PDEs that draws inductive bias from boundary-element methods and equivariant ML. GLOBE represents solutions as superpositions of learnable Green’s-function-like kernels evaluated from boundary faces to targets, composed across multiscale branches and communication hyperlayers. The architecture is translation-, rotation-, and parity-equivariant; discretization-invariant in the fine-mesh limit; and units-invariant via rigorous nondimensionalization. An explicit far-field decay envelope stabilizes extrapolation, boundary-to-boundary hyperlayer communication mediates long-range coupling, and the all-to-all boundary-to-target evaluation yields a global receptive field that respects PDE information flow, even for elliptic PDEs.
On AirFRANS (steady incompressible RANS over NACA airfoils), GLOBE achieves substantial accuracy improvements. On the “Full” split, it reduces mean-squared error by roughly $200\times$ on all fields relative to the dataset’s reference baselines, and roughly $50\times$ relative to the next-best-performing model. In the “Scarce” split, it achieves over $100\times$ lower error on velocity and pressure fields and over $600\times$ lower error on surface pressure than Transolver. Qualitative results show sharp near-wall gradients, coherent wakes, and limited errors under modest extrapolation in Reynolds number and angle of attack.
In addition to this accuracy, the model is quite compact (117k parameters), and fields can be evaluated at arbitrary points during inference. We also demonstrate the ability to train and predict with non-watertight meshes, which has strong practical implications.
These results show that rigorous physics- and domain-inspired inductive biases can achieve large gains in accuracy, generalizability, and practicality for ML-based PDE surrogates for industrial computer-aided engineering (CAE).
1 Introduction
1.1 Motivation: The Computational Challenge of CFD and the Need for Surrogates
Computational fluid dynamics (CFD) simulations are essential throughout modern engineering, from aerospace and automotive design to energy systems and climate modeling. In aerospace and automotive applications, the vast majority of engineering quantities of interest (e.g., time-averaged lift, drag, pressure distributions) are steady-state. Because of this, flows are commonly modeled using either a) Reynolds-averaged Navier-Stokes (RANS) equations or b) higher-fidelity (and inherently-unsteady) large-eddy simulation (LES) methods, from which a time-averaged solution is extracted. In either case, severe computational limitations arise for several reasons. First, the PDE’s nonlinearity makes iterative methods most practical [^1] Referring to outer-loop iterations for PDE linearization, which wrap any inner-loop iterations that are used to solve the linearized system itself., yet these introduce the possibility of convergence failures and unfavorable speed-stability tradeoffs. Second, the PDE’s elliptic character demands global information propagation and solution relaxation (discussed in Section 1.4), driving high computational cost – even with acceleration techniques such as multigrid or tailored initialization strategies (e.g., Sharpe et al. (2025)). Finally, at the high Reynolds numbers typical of industry interest, thin boundary layers develop on no-slip walls, requiring many degrees of freedom to resolve these high-gradient regions that are critical for accurate force predictions. Due to the combination of these factors, industrial CFD simulations routinely consume days to weeks of supercomputer time.
This computational expense has motivated rapid growth in machine-learning-based surrogate models that learn to approximate PDE solution operators from data. However, as discussed in Section 1.4, many existing ML surrogates do not structurally enforce fundamental physical requirements: they may break under coordinate transformations (lacking translation, rotation, or parity equivariance), depend on arbitrary discretization choices, lack unit consistency, or employ local receptive fields that limit generalization for elliptic PDEs. Architecturally enforcing these requirements can improve generalization beyond narrow training distributions, which is critical for industrial deployment.
1.2 Motivation from Boundary-Element Methods
Advanced numerical methods for aerodynamics provide instructive architectural guidance. Integral boundary layer (IBL) methods such as XFOIL by Drela (1989) achieve 100- to 1,000-fold computational speedups over volumetric RANS by discretizing only boundaries via boundary-element methods (BEM), exploiting the linear, homogeneous structure of potential flow. This boundary-centric formulation with global information transfer through boundary-only integrals encodes powerful inductive biases for fluid dynamics, and indeed more broadly to many PDEs as shown by Martinsson (2020). In this work, we combine such domain-inspired architectural ideas with equivariant ML and the flexibility of learned representations. The long-term goal is to extend the accuracy and speed benefits of IBL methods to complex 3D cases where analytical IBL extensions have long remained elusive. Appendix A provides additional background on IBL methods and their relevance to the GLOBE architecture introduced in Section 2.
1.3 Related Work: ML Surrogate Models for PDEs
Recent ML surrogate approaches offer different tradeoffs in handling unstructured meshes, discretization invariance, and geometric symmetries. Table 1 summarizes key methods [^2] A comprehensive review is available in Catalani et al. (2023).. CNNs excel on structured grids but struggle with unstructured meshes. GNNs handle irregular domains but require many layers for global coupling in elliptic PDEs. Neural operators achieve discretization invariance but often use localized kernels. Neural fields (Serrano et al., 2023; Catalani et al., 2025b) are closest to GLOBE in evaluating at arbitrary points, but typically compress boundary and problem information into fixed-size global latents. GLOBE differs by maintaining per-face latents, using kernel functions from each face rather than a single field, and structurally enforcing Euclidean equivariance and representational invariances throughout.
Table 1: ML surrogate approaches for PDEs and their characteristics.
Method Key Strengths Key Limitations
CNNs / UNets Hierarchical features; efficient on structured grids Structured grids only; lack geometric symmetries; costly interpolation for unstructured meshes
GNNs Handle unstructured meshes; effective for parabolic PDEs Local receptive fields require $\mathcal{O}(\text{diameter})$ layers for elliptic PDEs; typically lack rotation equivariance
Neural Operators Discretization invariance; function-space mappings; well-suited for heterogeneous PDEs that are chiefly driven by field forcing functions Often use localized kernels to incorporate geometry information, limiting information flow; limited ability to incorporate boundary information for boundary-driven problems
Neural Fields Arbitrary query points; continuous representations; discretization invariance Global compression bottleneck; limited geometric equivariance
1.4 Necessary Attributes of Physics Models
To motivate the architectural choices presented later in this work, we review several fundamental qualities that physics models – whether numerical solvers or learned surrogates – must satisfy. We organize these requirements into three categories: physical conservation laws, representation invariances, and problem-specific constraints. Table 2 provides a summary.
Table 2: Necessary attributes for valid physics models and their architectural implications.
Attribute Physical Basis Architectural Implication
Physical Conservation Laws & Symmetries
Translation-equivariance Conservation of momentum Use only relative spatial information; no absolute coordinates
Rotation-equivariance Conservation of angular momentum Vectors via local basis reprojection, not channel concatenation
Parity-equivariance Euclidean symmetry group Proper handling of vectors vs. pseudovectors
Representation Invariances
Discretization-invariance Numerical consistency Predictions must converge as mesh is refined
Units-invariance Physical consistency Operate in unitless space via nondimensionalization
Problem-Specific Constraints (examples for PDEs)
Receptive field structure PDE well-posedness Global receptive field for elliptic PDEs; match characteristic speeds for hyperbolic PDEs
Boundary condition encoding PDE well-posedness Explicit BC type labels as inputs
1.4.1 Physical Conservation Laws and Symmetries
These requirements follow directly from fundamental physics via Noether’s theorem, which connects conservation laws to continuous symmetries.
Translation-equivariance.
Any proposed physics model must obey conservation of momentum, and by Noether’s theorem, this requires translational symmetry: if all inputs are translated in space, outputs should be identically translated. Architecturally, this requires using only relative spatial information – any use of absolute coordinates in the core learned encoding of a model breaks this symmetry.
Rotation-equivariance.
Conservation of angular momentum requires rotational symmetry. While rare in ML surrogates for PDEs, rotation-equivariant architectures have achieved remarkable successes in other domains (Thomas et al., 2018; Fuchs et al., 2020; Jumper et al., 2021; Batzner et al., 2022), with libraries such as e3nn (Geiger & Smidt, 2022) enabling implementation. Common sources of violations include treating vector components as concatenated scalar channels (which implicitly fixes an arbitrary global Cartesian frame) or using traditional convolutional layers on structured grids. If a trained model does not produce identical output when the grid is rotated, it lacks rotational equivariance.
Parity-equivariance (reflection / chirality).
Almost all physical laws are equivariant under parity [^3] With the fascinating exception of the weak force at subatomic scales, as discovered by Wu et al. (1957).. Intuitively, a pair of vortices spinning clockwise and counterclockwise must have exactly identical decay rates. Combining parity symmetry with translation- and rotation-equivariance forms the Euclidean symmetry group. Structurally embedding Euclidean symmetry substantially improves data efficiency and generalizability (Geiger & Smidt, 2022; Batzner et al., 2022; Zhang, 2022; Sharpe & Hansman, 2025; Geiger, 2025). While equivariant ML research has historically focused on materials science, chemistry, and biology, these symmetries are equally relevant in computer-aided engineering (CAE) applications - yet their incorporation into PDE surrogates in this domain remains rare.
On the infeasibility of data augmentation for equivariance learning.
[…]
1.4.2 Representation Invariances
In addition to respecting physical symmetries, any valid physics model must produce results that are independent of arbitrary choices in how we represent and discretize the problem.
Discretization-invariance (numerical consistency).
For a valid solver, discretization error must approach zero as mesh resolution increases – this is numerical consistency in CFD. Architecturally, this requires that the model’s predictions converge to a well-defined limit independent of the specific mesh used. CNNs and GNNs tie their representations to specific discretizations; neural operators and neural fields achieve this property by parameterizing continuous functions.
Units-invariance.
Another fundamental requirement is invariance to the choice of physical units. The same physical scenario, whether represented in SI or Imperial units, should yield the same physical solution. This necessitates dimensional consistency throughout the entire modeling pipeline. Notably, most ML surrogates violate this principle during data preprocessing, before the model architecture is even considered – for example, by concatenating features with incompatible units (such as combining velocities, $L][T]^{-1}$, with lengths, $[L]$). A practical way to achieve units-invariance is to learn in unitless spaces via rigorous nondimensionalization, as in Sharpe & Hansman (2025).
1.4.3 Problem-Specific Constraints
These requirements vary depending on the PDE type and problem formulation, but are equally critical for learning valid solution operators.
PDE information flow and receptive fields.
Beyond the symmetries discussed above, any valid physics model must respect the information flow structure dictated by the PDE itself. Different PDE types [^5] Here, we restrict this classification discussion to second-order-and-higher PDEs, as these form the vast majority of computationally-expensive PDEs used in engineering applications. impose fundamentally different requirements on how information propagates through space:
Elliptic PDEs (e.g., steady-state aerodynamics, electrostatics, static structural analysis) exhibit global coupling: every point in the domain is influenced by every boundary condition and source term. Mathematically, well-posedness requires that boundary data at any location can affect the solution at any other location. For a surrogate model, this necessitates a global receptive field: the model’s output at any query point must have access to information from the entire domain.
Parabolic PDEs (e.g., heat diffusion, viscous flows) exhibit infinitely fast spatial information propagation, but distant influences decay rapidly with time. For small time steps $\Delta t$, local receptive fields can provide reasonable approximations; however, large-timestep solution operators still require spatially-distant information access.
Hyperbolic PDEs (e.g., wave propagation, inviscid compressible flow) have finite characteristic speeds. Information travels along characteristic curves at well-defined wave speeds. A valid solver must resolve these characteristics: the receptive field at time $t$ must encompass all points reachable via characteristics originating from initial/boundary data. This requirement is in some ways an analogue of the Courant-Friedrichs-Lewy (CFL) condition in explicit time-stepping: if the numerical receptive field fails to capture physical wave propagation, the solver becomes fundamentally unstable or inaccurate.
Architectures with local or limited-range receptive fields cannot satisfy well-posedness requirements for PDEs requiring global coupling. Vanilla GNNs, for instance, would need layers proportional to mesh diameter for elliptic PDEs - intractable to satisfy on arbitrarily-fine meshes. Many neural operator implementations similarly employ localized kernels (ball queries, k-NN) that lack global coupling.
When such architectures nonetheless achieve strong benchmark performance on elliptic PDEs despite these architectural limitations, it suggests they may be learning correlations with problem-specific features (mesh patterns, geometric heuristics) rather than the underlying PDE physics. For example, a local GNN applied to aerodynamics might learn that fine mesh cells correlate with boundary layer regions and hence lower velocities, or that distance to a boundary (measured by neighbor-hops) correlates with pressure gradients. Such patterns can produce good metrics on validation sets from the same distribution but fail catastrophically under genuine distribution shift to different geometries, flow regimes, or mesh strategies. For applications requiring robust out-of-distribution generalization, architecturally ensuring that receptive fields match the PDE’s intrinsic information flow provides stronger guarantees than relying on training data to implicitly capture these patterns.
Explicit boundary condition encoding.
A related and equally fundamental issue is that many PDE problem encodings in the ML literature fail to encode any boundary condition (BC) type information whatsoever, rendering the learning task mathematically ill-posed: it is not possible to truly learn a PDE’s solution operator if the boundary conditions are not provided. Consider an automotive aerodynamics case with mixed BCs: no-slip (vehicle body) and free-slip (ground plane). Without explicit BC labels, models either a) cannot distinguish boundary condition differences, leading to inaccurate solutions, or b) worse, they infer boundary condition types from nearby geometry, learning spurious correlations (e.g., “flat surfaces tend to be road surfaces, which tend to be slip walls”) that also fail catastrophically under distribution shift. Proper generalization requires BC types as explicit inputs, ensuring physically grounded learning rather than dataset artifacts.
1.4.4 Architecture Design Philosophy for ML Surrogates for PDEs
The prevalence of architectures in the field of ML for PDEs that violate these requirements reflects a mismatch between the domains where modern ML has historically achieved its greatest successes and the requirements of modeling physics-governed systems. While breakthroughs in perception, language, and game-playing have naturally influenced and advanced PDE surrogates, these domains differ fundamentally from physics in their treatment of structure and generalization. For example, symmetries can be useful heuristics in computer vision [^6] For example, see the data augmentation techniques in Krizhevsky et al. (2017), but the inevitable exceptions and edge cases limit the usefulness of structurally-enforcing these symmetries.
By contrast, in physics, conservation laws and coordinate invariances are exact, universal mathematical constraints. Hence, by encoding these symmetries, we reduce the hypothesis space to physically-plausible functions, improving data efficiency and out-of-distribution accuracy. In CAE applications, this out-of-distribution generalization is particularly crucial, due to adverse selection bias: engineers reach for new simulations precisely when encountering scenarios that are far from any previously-analyzed case. ML benchmarks with i.i.d. train/test splits may therefore inadvertently reward architectures that achieve high in-distribution accuracy through mechanisms that preclude practical generalization (encoding absolute position, architecture-specific discretizations, local receptive fields), as discussed by McGreivy & Hakim (2024).
As the field matures toward foundation-model aspirations, we believe that incorporating physical structure at the architectural level is essential. Foundation models must generalize across vast ranges of geometries, flow regimes, and boundary conditions. Models that violate conservation laws, break under coordinate transformations, or depend on arbitrary discretization choices are fundamentally learning correlations in the training distribution rather than underlying physics. If a model violates basic physical requirements known to be exact, it has likely learned dataset-specific patterns rather than a general solution operator, regardless of its performance on held-out test sets from the same distribution.
1.5 Contributions and Paper Organization
This paper’s primary contributions are:
  • A first-principles review of the necessary attributes that any valid physics model must possess, and guidance for constructing architectures that respect these (Section 1.4).
  • A new domain-inspired ML architecture for PDE surrogate modeling that respects these attributes, GLOBE (Section 2).
  • Evaluation of the GLOBE model on an airfoil aerodynamics case study (Section 3), including results, discussion, and comparison to state-of-the-art surrogate models (Section 4).
Novel technical contributions include:
  • Equivariant kernel functions that maintain exact translation, rotation, and parity symmetries through invariant feature engineering and local basis reprojection.
  • Padé-approximant MLPs whose rational-function structure naturally matches Green’s function decay characteristics, providing stronger inductive bias for physics kernels.
  • Multiscale kernel composition that explicitly encodes multiple characteristic length scales (viscous, geometric, and their learned combinations), enabling efficient representation across disparate spatial scales.
  • Communication hyperlayers that propagate boundary condition information between BC partitions before interior evaluation, learning source strengths in a manner inspired by but computationally lighter than traditional boundary-element linear solves.
  • Rigorous nondimensionalization throughout the data pipeline, ensuring units-invariance and physical consistency.
Section 2 details the GLOBE architecture; Sections 3 and 4 demonstrate its application to aerodynamic surrogate modeling.
2 GLOBE Model Architecture
We now introduce GLOBE: Green’s-function-Like Operator for Boundary Element PDEs. The architecture synthesizes three key insights: (1) for linear homogeneous PDEs, solutions can be exactly represented as superpositions of Green’s functions; for nonlinear PDEs, learnable kernel superpositions can approximate solutions when the PDE operator is weakly nonlinear or exhibits approximately linear behavior in significant domain regions; (2) for boundary-driven problems, discretizing only boundaries (not interiors) dramatically reduces degrees of freedom while maintaining global information flow; and (3) neural networks can learn kernel shapes that generalize Green’s function concepts to settings where analytical forms are unknown or do not exist.
While the theoretical framework is general, in this work we develop and evaluate GLOBE specifically for aerodynamic flows, where the RANS equations admit quasi-linear approximations. Specifically, at high Reynolds numbers and with mostly-attached flows, thin shear layer approximations apply, as described in full detail by Drela (2013a; 2019). This allows decomposition of the flow into potential flow (linear) and boundary layers. The latter are governed by the von Kármán momentum integral equation, which is parabolic in the streamwise coordinate and hence far more tractable than the full RANS system; furthermore, the surface-normal dimension can be simplified away via integral boundary layer theory, as in Drela (1989). This motivates kernel-based representations even for nominally nonlinear PDEs. The quasi-linear tractability of RANS under industrially-relevant assumptions is particularly significant: if even this notoriously nonlinear system admits effective kernel-based treatment, the approach likely extends to other PDEs exhibiting regional or weak nonlinearity. More broadly, the architecture is designed for PDEs with the following characteristics:
  • Homogeneous governing equations: no spatially-varying forcing terms or material properties
  • Boundary-driven physics and focus: quantities of interest are primarily driven by and/or measured at boundaries (e.g., boundary layers, stress concentrations)
  • Elliptic or mixed character: requiring global information propagation for a solver to be well-posed
  • Weak or regional nonlinearity: PDEs where the governing operator is linear or exhibits approximately linear behavior in significant portions of the domain
  • Multiscale structure: PDEs exhibiting diverse phenomena at multiple characteristic length scales
This profile encompasses many industrially critical problems in computational engineering. We focus our evaluation on external aerodynamics, though the architectural principles may extend to other boundary-driven elliptic problems such as static structural analysis, heat conduction, and electrostatics. Demonstrating such transferability remains important future work.
GLOBE learns a mapping shown in Figure 1. This mapping is similar to those found in neural-field-based architectures, and GLOBE could perhaps be classified as a kernel-based neural field variant. However, the GLOBE model architecture has crucial differences from traditional neural fields that (a) allow it to respect the necessary attributes of a valid physics model, such as Euclidean equivariance, and (b) avoid the global compression bottleneck common in neural fields. While GLOBE does propagate latent variables through communication hyperlayers, these latents are physically bound to individual boundary faces rather than compressed into a single global vector.
More generally, GLOBE actually has more in common with traditional boundary-element-method PDE solvers than it does with existing popular ML surrogate architectures. In particular, as discussed in Section 1.2, heavy inspiration is drawn from aerodynamic panel methods and integral boundary layer methods such as those described by Katz & Plotkin (2004); Drela (1989; 2013a; 2019); Zhang (2022).
Problem Inputs A sequence of one or more boundary meshes (one per BC type), each of which consists of: $\circ$ Geometry: triangle faces in 3D or line edges in 2D; represented as vertices and face indices (e.g., an unstructured STL file) $\circ$ The boundary condition (BC) type as a string label (e.g., “no_slip”, “inlet”, “outlet”) $\circ$ Any unitless scalar or vector values associated with this BC type (e.g., values of a Dirichlet boundary condition) A set of scalar, dimensional reference lengths $\{\ell_{i}\}$, obtained through combinations of dimensional parameters in the governing equations (via Buckingham $\pi$ theorem or domain expertise) (Optional) Global scalars and/or vectors (unitless), which could be drawn from the governing equations or infinite far-field boundary conditions Query Points Spatial coordinates where the solution fields are evaluated GLOBE Model (Figure 2) Output Values of the solution fields (scalars and/or vectors, all unitless) at the query points
Figure 1: The input-output mapping of a GLOBE model.
2.1 Architecture Overview
GLOBE predicts fields on arbitrary query points in space, given one or more boundary meshes partitioned by boundary-condition (BC) type. The architecture is organized in three levels, most easily explained bottom-to-top:
  • At the lowest level, the centroid of each face of a boundary mesh acts as a source point that influences every target point in an all-to-all manner via a learned kernel function $\mathcal{K}$. This kernel is shared across all faces. Most of the desirable physical properties of the overall model (e.g., equivariances) are derived through the careful structure of this component, so this is the most critical layer. Further description is provided in Section 2.2.
  • At the middle level, GLOBE implements a multiscale composition of multiple kernel functions, whose influences are linearly superimposed. Each individual kernel is still shared across boundary faces, and is associated with a different reference length scale $\ell_{i}$. These length scales are derived from combinations of dimensional constants within the PDE governing equations, and should be given as model inputs. This multiscale composition is described in Section 2.3.
  • At the top level, GLOBE implements a stack of communication hyperlayers, which is a sequence of these multiscale kernel compositions. These hyperlayers pass latent scalar and vector fields from all boundaries to all other boundaries, prior to the final evaluation at user-chosen query points. These hyperlayers are also used to compute the source strength values associated with each face, which are used to linearly weight the contributions of each source face to each target point in the subsequent hyperlayer. This is described in Section 2.4.
Figure 2 depicts the overall architecture. Boundary meshes are merged by BC type; within each hyperlayer, the kernel is evaluated from every source BC to targets taken from either (a) other boundary face centers (to propagate latent variables) or (b) user query points (final layer). Outputs across source BC types are summed. After the final prediction, a learnable calibration (affine for scalar fields, linear for vector fields777required to preserve rotation-equivariance) is applied.
Top Level: Communication Hyperlayers (Section 2.4) Inputs Boundary meshes; Reference lengths; Global scalars & vectors Query points Initialize $w_{s}^{(k,0)}=1$ latents $\emptyset$ Hyperlayer 0 $\sum_{i}$ Multiscale kernels, with: Sources: $\text{BC}_{i}$ faces Targets: all faces Outputs: $w_{s}^{(k)}$ for $H_{1}$ kernels; latent scalars & vectors Hyperlayer 1 $\sum_{i}$ Multiscale kernels, with: Sources: $\text{BC}_{i}$ faces Targets: all faces Outputs: $w_{s}^{(k)}$ for $H_{2}$ kernels; latent scalars & vectors $\cdots$ Hyperlayer $H$ $\sum_{i}$ Multiscale kernels, with: Sources: $\text{BC}_{i}$ faces Targets: query pts. Outputs: raw solution fields at query points Calibration Scalars: affine; Vectors: linear Output Solution fields at query pts. latent data Zooming into a single multiscale kernelassociated with a single source BC typeMiddle Level: Multiscale Kernel (one per source BC type, e.g., ‘inlet’, ‘outlet’, ‘wall’; Section 2.3) Inputs Source points $\{\bm{x}_{s}\}$; Target points $\{\bm{x}_{t}\}$; Source scalars & vectors; Global scalars & vectors; Source strengths $w_{s}^{(k)}$ (per branch); Log-ratios $\ln(\ell_{i}/\ell_{j})$ Kernel $\mathcal{K}^{(1)}$ with ref. length $\ell^{\text{eff}}_{1}=\ell_{1}\,\exp(\alpha_{1})$ Kernel $\mathcal{K}^{(2)}$ with ref. length $\ell^{\text{eff}}_{2}=\ell_{2}\,\exp(\alpha_{2})$ $\vdots$($n$ branches) Kernel $\mathcal{K}^{(n)}$ with ref. length $\ell^{\text{eff}}_{n}=\ell_{n}\,\exp(\alpha_{n})$ $\sum_{k}$ Output Field values at targets $\bm{u}_{t}$ (Eq. 9) Each kernel function has its own learnable parameters; shared architecture (Sec. 2.3) Zooming into a single kernel $\mathcal{K}^{(k)}$Bottom Level: Kernel Function $\mathcal{K}$ (source $\to$ target evaluation; Section 2.2) Input Assembly Relative position $\bm{r}_{ts}=\frac{\bm{x}_{t}-\bm{x}_{s}}{\ell}$; Source scalars $\bm{s}$; Source vectors $\bm{V}$; Global scalars & vectors Invariant Features Scalars $\bm{s}$; $\operatorname{smoothlog}(\left\lVert\bm{v}_{i}\right\rVert)$ for vectors; Weighted $\operatorname{P}_{i}(\cos\theta_{ab})$ on vector pairs Padé MLP Process concat. scalar features Far-Field Decay Scale by envelope (Eq. 6): $\frac{1-\exp(-\left\lVert\bm{r}\right\rVert^{2})}{(\left\lVert\bm{r}\right\rVert^{2}+1)^{(d-1)/2}}$ Vector Reprojection Project onto basis $\bm{B}=[\bm{\hat{r}},\bm{\hat{v}}_{i},\ldots]$ Aggregation $\sum_{s}w_{s}\,a_{s}\,\mathcal{K}_{ts}$ (Eq. 7) Output Aggregated kernel output $\mathcal{K}_{t}$ Shared kernel parameters across all $(t,s)$ pairs; aggregate over sources (Eq. 7) Equivariance Mechanisms:
Translation: relative positions only Rotation: invariant features + reprojection Parity: parity-preserving reprojection Discretization: area-weighted aggregation Units: rigorous nondimensionalization
Figure 2: GLOBE architecture. The model operates at three hierarchical levels. Top: Communication hyperlayers propagate latent information between boundary partitions before final evaluation. Middle: Each hyperlayer contains multiscale kernels operating at different reference length scales, with outputs summed. Bottom: Individual kernels evaluate all-to-all source-to-target influences through invariant feature engineering, Padé-approximant networks, far-field decay envelopes, and equivariant vector reprojection, then aggregate over sources weighted by strengths and face areas.
Mathematical properties.
The architecture satisfies the necessary attributes established in Section 1.4: translation-, rotation-, and parity-equivariance through relative positions and local reprojection onto carefully-constructed vector bases; discretization-invariance via area-weighted boundary integrals; and units-invariance through rigorous nondimensionalization. More detailed discussion of these properties is provided in Appendix D. Additionally, GLOBE’s all-to-all boundary-to-target evaluation (Eq. 7) provides a true global receptive field: every query point receives information from every boundary face, directly fulfilling elliptic PDE requirements. For parabolic and hyperbolic PDEs, global receptive fields remain beneficial (allowing the model to learn long-range couplings if present) without imposing structural limitations. Communication hyperlayers (Section 2.4) further enable boundary-to-boundary information exchange, mediating the kind of global boundary coupling that traditional boundary-element methods achieve through dense linear solves.
2.2 Kernel Functions
A GLOBE model fundamentally represents the solution of a PDE as a linear combination of learnable kernel functions $\mathcal{K}$. These kernels were designed with inductive biases inspired by Green’s functions: in the case of a linear PDE, the ideal kernels would be the PDE’s Green’s functions themselves. We provide visual illustrations of the kernel’s properties and symmetries in Appendix B to aid intuition before proceeding with the mathematical development.
The kernel function $\mathcal{K}$ computes influence values for each output field (whether scalar or vector) from each source face to each target point. For simplicity, in the following subsections we consider the base case of a single source face’s influence on a single target point. We first describe features, then the network backend and physics constraints, and finally aggregation.
2.2.1 Inputs
We begin by gathering all relevant scalars and vectors associated with this particular source-to-target interaction. We denote this collection of scalars as $\bm{\xi}=[\xi_{1},\xi_{2},\ldots,\xi_{n}]$ and vectors as $\bm{V}=[\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{m}]^{T}$.
To start, we take a set of scalar- and vector-valued inputs to the kernel function as elements of $\bm{\xi}$ and $\bm{V}$, respectively. Some of these may be associated with the source face (e.g., boundary condition values or latent variables from prior processing); others may be global (e.g., parameters of the governing PDE). Critically, all inputs should be physically dimensionless to preserve units-invariance.
In addition to any such scalar and vector quantities that are given as inputs, two are added as engineered features: the unit normal vector of the source face $\bm{\hat{n}}_{s}$, and the relative position vector from the source face center to the target point $\bm{r}_{ts}$. This latter vector must be nondimensionalized by a reference length scale $\ell$ to preserve units-invariance:
$\bm{r}_{ts}\,=\,\frac{\bm{x}_{t}-\bm{x}_{s}}{\ell}\in\mathbb{R}^{d},$ (1)
  • $\bm{x}_{t}$ spatial position of target point $t$
  • $\bm{x}_{s}$ spatial position of source face center $s$
  • $\ell$ reference length scale associated with each kernel branch; provided as input to the multiscale composition (Section 2.3)
  • $d$ number of spatial dimensions (typically 2 or 3)
Because these raw inputs contain only relative positional information from PDE boundaries to target points, the kernel function itself becomes translation-equivariant.
2.2.2 Invariant Engineered Features
The scalar collection $\bm{\xi}$ can be used directly as model inputs. However, the vector collection $\bm{V}$ cannot be used directly without breaking rotational-equivariance. We therefore transform these vectors into rotationally-invariant features. (Rotational equivariance is later restored by vector reprojection in Section 2.2.5.)
One possible rotationally-invariant feature is the magnitude $\left\lVert\bm{v}_{i}\right\rVert$. However, this has two issues: (a) very large magnitudes can dominate the scalar collection, saturating downstream steps, and (b) the magnitude function breaks $C^{1}$ continuity near the zero vector, violating continuous differentiability [^8] While not strictly required for surrogate modeling, many practical use cases (e.g., design optimization) benefit heavily from this property..
Instead, we encode the value
$\operatorname{smoothlog}(\left\lVert\bm{v}_{i}\right\rVert)$ (2)
for each vector $\bm{v}_{i}$ as an additional scalar in the collection $\bm{\xi}$. Here, we define the function $\operatorname{smoothlog}:\mathbb{R}^{1}\to\mathbb{R}^{1}$ as
$\operatorname{smoothlog}(x)=(1-e^{-x})\cdot\ln(1+x),$ (3)
which is crafted to have several desirable properties on the domain $x\geq 0$ (as is the case when applied to the raw magnitude values):
  • $\operatorname{smoothlog}(x)\sim x^{2}$ as $x\to 0$
  • $\operatorname{smoothlog}(x)\sim\ln(x)$ as $x\to\infty$
  • $\operatorname{smoothlog}(x)$ is monotonically increasing for $x>0$, and hence forms a bijection with $x$
  • $\operatorname{smoothlog}(x)$ is $C^{\infty}$ continuous for all $x\geq 0$, with higher-degree derivatives that grow relatively slowly with degree
  • $\operatorname{smoothlog}(x)$ can be efficiently and accurately computed using the expm1 and log1p special functions
Figure 3 illustrates these asymptotic properties.
Figure 3: The $\operatorname{smoothlog}(x)$ function and its derivative, with near- and far-field asymptotic limits shown as dashed lines.
Next, we compute additional engineered features by considering every possible unordered pair of vectors $\bm{v}_{a}$ and $\bm{v}_{b}$ in the collection $\bm{V}$. For each such pair, we evaluate the magnitude-weighted spherical harmonic features
$\operatorname{smoothlog}\!\left(\left\lVert\bm{v}_{a}\right\rVert\cdot\left\lVert\bm{v}_{b}\right\rVert\right)\cdot\operatorname{P}_{i}\!\left(\cos(\theta_{ab})\right),$ (4)
where $\operatorname{P}_{i}$ is the $i$-th Legendre polynomial (for $i=1,2,\ldots,n_{\text{harmonics}}$, a hyperparameter), and $\theta_{ab}$ is the angle between vectors $\bm{v}_{a}$ and $\bm{v}_{b}$. These are added to the collection $\bm{\xi}$. Multiplication by $\operatorname{smoothlog}\!\left(\left\lVert\bm{v}_{a}\right\rVert\cdot\left\lVert\bm{v}_{b}\right\rVert\right)$ is necessary to preserve $C^{\infty}$ continuity: without this scaling, $\theta_{ab}$ can change discontinuously as either vector crosses zero. Effectively, Eq. 4 forms a special case of a spherical harmonics expansion on the pairwise angle, weighted by a nonlinear function of the magnitudes.
At this point, all relevant information has been gathered and encoded into a collection of scalars $\bm{\xi}$. As scalar-valued quantities, these are inherently rotationally-invariant (in addition to translation-invariant, from Section 2.2.1), facilitating further processing.
2.2.3 Processing with Padé-Approximant MLPs
At this stage, the scalar collection $\bm{\xi}$ could be processed by a standard multilayer perceptron (MLP). However, we find empirically that using a custom Padé-approximant MLP instead yields substantial improvements in model accuracy and training speed.
We introduce what we call a “Padé-approximant MLP” as a learnable mapping of the form:
$\text{Padé}\left(\bm{\xi}\right)\;=\;\frac{\operatorname{sgn}\!\left(\phi_{n}\left(\bm{\xi}\right)\right)\,\left\lvert\phi_{n}\left(\bm{\xi}\right)\right\rvert^{N}}{1+\left\lvert\phi_{d}\left(\bm{\xi}\right)\right\rvert^{D}}.$ (5)
Here, $\phi_{n}$ and $\phi_{d}$ are independent learnable MLPs with $C^{\infty}$ differentiable activations (SiLU in the reference implementation); $N$ and $D$ are the integer-valued orders of the numerator and denominator, respectively. The numerator uses sign-preserving exponentiation, while the denominator uses magnitude-only exponentiation; when combined with the constant term, this ensures that the denominator is always $\geq 1$, preventing arbitrarily large Lipschitz constants.
We typically set $N=2$ and $D=2$; while other choices can be made, the property $N=D$ is quite useful as it ensures that the Padé-approximant layer asymptotes to a constant value in the far-field limit, for any direction in the input space $\bm{\xi}$. This will later aid in careful control of the asymptotic behavior of the kernel function, as discussed in Section 2.2.4.
The motivation stems from Green’s function structure: these typically exhibit rational-function-like behavior with algebraic singularities near sources and algebraic decay at infinity (e.g., $1/r$ for 3D Poisson). While standard MLPs are universal approximators, Padé-approximant MLPs provide explicit inductive bias toward this functional form, enabling more parameter-efficient learning of physics kernels. Appendix C provides visualizations demonstrating the expressivity of randomly-initialized Padé-approximant MLPs compared to standard MLPs.
2.2.4 Far-Field Decay Constraint
To bias the model toward physically plausible decay with distance, we then scale all the raw outputs of the Padé-approximant MLP by a function of the relative position vector $\bm{r}_{ts}$:
$g\big(\left\lVert\bm{r}_{ts}\right\rVert\big)\,=\,\frac{1-e^{-\left\lVert\bm{r}_{ts}\right\rVert^{2}}}{\left(\left\lVert\bm{r}_{ts}\right\rVert^{2}+1\right)^{(d-1)/2}},$ (6)
where $d$ is the number of spatial dimensions (2 or 3), yielding $\sim 1/r$ decay in 2D and $\sim 1/r^{2}$ in 3D as $\left\lVert\bm{r}_{ts}\right\rVert\to\infty$ in any direction.
This introduces two useful effects. First, self-influence at $\bm{r}_{ts}=\bm{0}$ is driven to zero, encouraging kernels to share influence across faces and preventing overly compact support (the “bed-of-nails interpolant” failure mode in radial basis interpolation). The chosen decay is inspired by the Lamb-Oseen vortex profile. Second, the far-field decay encodes an inductive bias toward the inverse square law: a common large-distance behavior of particle-mediated interactions (or alternatively, Gauss’s law for fields with conserved fluxes) that is observed across a vast array of physical systems, while still allowing the model to learn deviations where appropriate (e.g., advection-dominated regimes).
This also introduces favorable generalization properties in spatial extrapolation. When a GLOBE model is queried with target points that are much more distant from PDE boundaries than any target points seen in the training data, the kernel function cannot extrapolate wildly: “in the absence of other information, physical phenomena tend to zero at infinity.” Additionally, this guaranteed far-field decay enables the formulation of infinite far-field boundary conditions, where boundary effects naturally vanish at large distances without requiring explicit far-field boundary meshes.
2.2.5 Vector Reprojection for Rotational Equivariance
The Padé-approximant MLP outputs a latent vector $\bm{z}$ that is not a physical-space vector; its dimensionality typically far exceeds $d$. Interpreting any subset of channels as a physical vector would implicitly select an arbitrary global Cartesian frame, breaking rotational symmetry. Therefore, while individual channels of $\bm{z}$ can be directly interpreted as physical scalar fields, constructing physical vector outputs requires a separate procedure.
To illustrate how we can achieve this, consider the simplified case of constructing a single physical vector $\bm{u}$ associated with the pairwise interaction between a source face and a target point. We fundamentally need to project channel(s) of $\bm{z}$ onto a local basis vector set with non-arbitrary directionality to preserve rotational symmetry. A convenient source of non-arbitrary directionality can be found in the set of vectors $\bm{V}$ that were gathered as inputs to the kernel function (in Section 2.2.1). We then assemble a set of basis vectors [^9] Strictly speaking, this is not a true basis in the linear algebra sense, as (a) we typically have more basis vectors than spatial dimensions, and (b) the collection may not always span all of $\mathbb{R}^{d}$ (e.g., if every vector in $\bm{V}$ is the zero vector). This is not problematic for our approach, but it is worth noting that such a reprojection is generally not invertible. $\bm{B}=[\bm{b}_{1},\bm{b}_{2},\ldots,\bm{b}_{n}]$ by including the following components:
  1. 1.
    Base direction:
    • The radial unit vector $\bm{\hat{r}}$ from the source face to the target point.
  2. 2.
    For each non-$\bm{r}$ input vector $\bm{v}_{i}$ with unit direction $\bm{\hat{v}}_{i}$:
    1. (a)
      Add the axis direction $\bm{\hat{v}}_{i}$.
    2. (b)
      In the 2D case ($d=2$): include a dipole-like direction $\bm{e}_{\kappa}$ (projection of $\bm{\hat{v}}_{i}$ onto the line orthogonal to $\bm{\hat{r}}$).
    3. (c)
      In the 3D case ($d=3$): include a polar/meridional direction $\bm{e}_{\theta}$ (projection of $\bm{\hat{v}}_{i}$ onto the plane orthogonal to $\bm{\hat{r}}$).
    4. (d)
      Finally, scale each of the above basis vectors (steps within this enumeration) by $\operatorname{smoothlog}(\left\lVert\bm{v}_{i}\right\rVert)$. Similar to the use of $\operatorname{smoothlog}(\cdot)$ in Section 2.2.2, this is subtly required to preserve $C^{\infty}$ continuity of the resulting field, and has the side benefit of mitigating saturation effects.
Optional tangential/vortex basis vectors. The above construction preserves parity symmetry. However, for physical phenomena that are strongly linked to vorticity, additional tangential basis vectors can provide useful inductive bias: in 2D, a tangential direction $\bm{e}_{\theta}$ orthogonal to $\bm{\hat{r}}$; in 3D, an azimuthal direction $\bm{e}_{\phi}$ orthogonal to both $\bm{\hat{r}}$ and $\bm{e}_{\theta}$. These directions are pseudovectors rather than true vectors, so including them naïvely breaks parity symmetry. A parity-preserving implementation is possible but requires explicit tracking of vector versus pseudovector semantics throughout the architecture, which is not yet implemented in the reference code. Consequently, these tangential/vortex-like directions are disabled by default, though users may enable them for applications where the inductive bias outweighs the loss of exact parity equivariance.
2.2.6 Aggregation over Sources
Let $\mathcal{K}_{ts}$ denote the kernel output value for the target-source pair $(t,s)$. The aggregated kernel value at each target point is a weighted sum over sources, implemented as a batched contraction:
$\mathcal{K}_{t}\,=\,\sum_{s=1}^{N_{s}}w_{s}\,a_{s}\,\mathcal{K}_{ts},$ (7)
  • $\mathcal{K}_{t}$ aggregated kernel output at target point $t$ (may be scalar or vector, depending on the output field)
  • $w_{s}$ source strength associated with each source face $s$; these are obtained at a higher level through the hyperlayer stack described in Section 2.4
  • $a_{s}$ area of the source face $s$; multiplication by this area term is critical, as it yields discretization-invariance in the limit case of a refined source mesh
  • $\mathcal{K}_{ts}$ kernel output value for the target-source pair $(t,s)$
The resulting fields are then returned as a collection of scalars and vectors, each with batch size $(N_{t})$.
2.2.7 Complexity and Scalability via Hierarchical Acceleration
The complete architecture with communication hyperlayers (Section 2.4) incurs $\mathcal{O}(N_{s}^{2}+N_{s}\times N_{t})$ complexity. During training, the boundary-to-boundary $\mathcal{O}(N_{s}^{2})$ term dominates due to the ability to downsample to a reasonably low number of query points. At inference, where we typically evaluate on a high number of query points to populate complete results, the latter boundary-to-target $\mathcal{O}(N_{s}\times N_{t})$ term typically dominates. In both cases, this quadratic scaling is reminiscent of transformer self-attention mechanisms, and indeed the kernel evaluation shares conceptual similarities with attention: both implement all-to-all communication to propagate contextual information, though GLOBE’s kernels omit the softmax normalization to preserve kernel linearity (a useful inductive bias on linear- or nearly-linear PDEs) and discretization-invariance.
While this quadratic scaling limits naive application to massive meshes, two promising paths forward mitigate this. First, the robustness to decimation later demonstrated in Section 4.3 provides an immediate solution: training can be performed on heavily downsampled boundary representations ($n_{\text{src, train}}\ll n_{\text{src, true}}$) while still retaining good accuracy.
In the longer-term, the guaranteed far-field decay from Section 2.2.4 enables hierarchical approximation schemes analogous to fast multipole methods (FMM) or Barnes-Hut algorithms. The key insight is that distant source clusters can be approximated by aggregations at their centroids, reducing far-field interactions from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$ with careful tree construction. We leave implementation of such acceleration to future work, though the architectural design already accommodates this due to the kernel’s smoothness (inherited from $C^{\infty}$ continuity of $\operatorname{smoothlog}$ and Padé layers) and decay.
Thus, the combination of decimation-robust training and hierarchical inference offers a clear path to scalability for industrial 3D problems. This contrasts with many neural field architectures whose general-purpose MLP representations provide minimal bounds on far-field behavior, precluding hierarchical methods entirely.
2.3 Multiscale Kernel Functions
Many physical systems exhibit multiple characteristic lengths. In GLOBE, each reference length $\ell_{i}$ defines a parallel kernel branch. These branches share the same kernel function architecture (as defined in Section 2.2) and inputs, but have independent learnable parameters and have their maximum expressivity at different effective reference length scales. Their outputs are then linearly summed. Collectively, we refer to this collection of kernel branches as a multiscale kernel.
More precisely, for each index $i$ in the user-provided set of reference lengths, a separate kernel branch $\mathcal{K}^{(i)}$ is instantiated and evaluated with its effective reference length
$\ell^{\text{eff}}_{i}\;=\;\ell_{i}\,\exp(\alpha_{i}),$ (8)
  • $\ell_{i}$ user-provided reference length for branch $i$
  • $\alpha_{i}$ learned scalar that fine-tunes the scale during training
Because the relative position feature $\bm{r}_{ts}=(\bm{x}_{t}-\bm{x}_{s})/\ell^{\text{eff}}_{i}$ is normalized by $\ell^{\text{eff}}_{i}$ (to restate Eq. 1), each branch effectively “zooms” the same geometric interaction to a different scale.
During multiscale evaluation, we also add global scalar features exposing scale relationships: for every pair $(\ell_{i},\ell_{j})$ of reference lengths, we append $\ln(\ell_{i}/\ell_{j})$ to the global scalar input [^10] With logic such that the assignment order to numerator and denominator is always consistent.. This provides explicit, unitless information about relative scales of problem physics. (Encoding individual length values directly would break units-invariance.) This design introduces a key inductive bias toward scale similarity: the network behaves equivariantly [^11] Strictly speaking, when encoding all reference lengths from Buckingham $\pi$ theorem applied to governing equation parameters. under absolute scale changes so long as all nondimensional parameters (e.g., Reynolds number, Mach number, etc.) are fixed.
Let $\mathcal{K}^{(k)}_{ts}$ denote the output for kernel branch $k$. With per-source, per-branch strengths $w^{(k)}_{s}$ (Section 2.4), the multiscale field is
$\bm{u}_{t}\;=\;\sum_{k}\;\sum_{s=1}^{N_{s}}w^{(k)}_{s}\,a_{s}\,\mathcal{K}^{(k)}_{ts}$. (9)
  • $\bm{u}_{t}$ final field value (may be scalar or vector, depending on the output field) at target point $t$
  • $k$ kernel branch index, corresponding to each reference length scale
  • $w^{(k)}_{s}$ source strength for source face $s$ in kernel branch $k$
  • $a_{s}$ area of the source face $s$
  • $\mathcal{K}^{(k)}_{ts}$ kernel output value for target-source pair $(t,s)$ from kernel branch $k$
This extends Eq. 7 from a single kernel to a superposition across multiple kernel branches indexed by $k$, where each branch operates at a different effective reference length scale.
In the reference implementation, evaluation of the multiscale kernel also supports optional chunking over target points to control memory [^12] At training time, chunking has no effect on memory due to the need to store intermediate primals on the compute graph; memory reduction is better achieved by downsampled training. However, at inference time, chunking can drive down the peak memory footprint to extremely low levels (700 MB at a chunk size of 128 query points), due to the independence of pairwise interactions. without changing results.
Choosing reference lengths.
The set $\{\ell_{i}\}$ is a model design choice made by the user, and should be guided by the governing physics. For maximum model generalizability, these should be fundamental physical length scales from the governing equations; however, problem-specific reference lengths could be included for models designed with a tighter scope of use. For example, in external aerodynamics applications, candidates for reference lengths might include:
  • The viscous length $\nu/U_{\infty}$, where $\nu$ is the kinematic viscosity and $U_{\infty}$ is the freestream velocity
  • A problem-specific characteristic length, such as an airfoil chord length $c_{\text{ref}}$
  • The trailing-edge boundary layer thickness under Blasius and/or turbulent models (e.g., $\delta_{\text{lam}}=5\,c_{\text{ref}}/\sqrt{\rm Re_{x}}$ and $\delta_{\text{turb}}\approx 0.37\,c_{\text{ref}}/\rm Re_{c}^{1/5}$, where $\rm Re_{c}=U_{\infty}c_{\text{ref}}/\nu$ is the Reynolds number based on a characteristic streamwise length), depending on the problem
  • In rarefied flows, the mean free path length $\lambda$
In practice, using a small set of $\{\ell_{i}\}$ (typically 2 to 4) is sufficient. Physical relevance of each length’s variation matters more than absolute values; the learnable factors $\alpha_{i}$ adapt initial scale choices during training.
2.4 Communication Hyperlayers
The top level stacks $H$ communication hyperlayers that exchange latent information between boundary-condition (BC) partitions in an all-to-all manner before the final interior evaluation. Each hyperlayer contains one multiscale kernel per source BC type: all mesh faces with a “no slip” BC share one multiscale kernel, while all “free slip” faces share another. This enables GLOBE to learn BC-type-specific behavior, such as boundary layer presence or absence. These hyperlayers also compute the source strengths $w_{s}$ and $w^{(k)}_{s}$ introduced in Eqs. 7 and 9.
Per-hyperlayer evaluation.
For hyperlayer $h_{i}$ ($i=0,\dots,H$), we perform:
  1. 1.
    Initialization (only for $i=0$). For each source BC type, set per-branch source strengths $w^{(k,0)}_{s}=\mathbf{1}$ for all reference-length branches $k$ and source faces $s$. Here, the superscript $(k,i)$ denotes kernel branch $k$ at hyperlayer $i$; this notation extends the per-branch strengths $w^{(k)}_{s}$ from Eq. 9 to track distinct values across hyperlayers. Initialize latent scalars/vectors as empty.
  2. 2.
    Inputs per source BC. Assemble kernel inputs from per-face data and, for $i>0$, include the latent scalars/vectors produced by the previous hyperlayer $h_{i-1}$. Always include the unit normal vector as an input (Section 2.2.1).
  3. 3.
    Targets. If $i<H$, choose the target points to be the boundary face centers (grouped by target BC). If $i=H$, instead choose the target points to be the user-requested query points (generally, but not necessarily, in the interior of the domain).
  4. 4.
    Multiscale evaluation and aggregation. For each source BC type, evaluate its multiscale kernel at the current targets using the effective reference lengths (Eq. 8). During source-summation (Section 2.2.6), weight contributions by source strengths and face areas.
  5. 5.
    Outputs. If $i<H$, output a collection of tensors for each target BC boundary mesh containing (a) per-branch strengths (one scalar channel per reference length), denoted $w^{(k,i+1)}_{s}$, and (b) a small set of latent scalar/vector channels; these become inputs to $h_{i+1}$. If $i=H$, instead output the named physical fields at the query points, usually in the interior of the domain.
  6. 6.
    Final calibration. After the last hyperlayer only ($i=H$), a lightweight per-field linear calibration is applied: scalars use an affine map (scale and bias), vectors use a scale-only map. This preserves rotational equivariance while allowing data-driven calibration of magnitudes and offsets.
These steps implement boundary-to-boundary message passing across $H$ hyperlayers, followed by a boundary-to-interior evaluation, enabling long-range coupling across BC partitions.
Inductive bias.
The architecture of the communication hyperlayers is inspired by the linear solve step of traditional boundary-element methods. There, the strengths $\bm{w}$ associated with each Green’s function kernel are obtained by solving a linear system of the form $\bm{A}\bm{w}=\bm{b}$, where $\bm{A}$ corresponds to influence coefficients between Green’s functions and is typically square and dense; $\bm{b}$ are the boundary condition residuals in the $\bm{w}=\bm{0}$ case. While future work could explore using this linear solve step, we choose to use this hyperlayer architecture instead for several reasons:
  • This linear solve would quickly dominate the computational cost of the overall model, particularly as the number of boundary faces grows - indeed, this is often seen with traditional boundary-element methods. This is particularly egregious in the case of a direct linear solve (which introduces $\mathcal{O}(N_{s}^{3})$ time complexity). If we instead use an iterative linear solve, the compute cost may be better; however, if we compute parameter gradients using typical automatic differentiation, the memory requirements explode due to the need to store intermediate iterations. We could bypass the memory issue with a custom vector-Jacobian product rule, but this (a) introduces significant implementation complexity and (b) yields gradients that are only approximate (since the truncation error in the forward iterative solve is not reflected in the backward pass).
  • This linear solve step requires symbolic knowledge of the governing PDEs (to evaluate boundary condition residuals). This is not necessarily a problem, though it does increase implementation complexity, especially since the need for symbolics has so far been avoided in this architecture.
  • The hyperlayer architecture allows information about boundary condition values to propagate from boundary to boundary, which is critical in some PDE scenarios. For illustrative purposes, consider an external aerodynamics problem which is formulated with an explicit “freestream” boundary mesh at some large finite distance to the body [^13] This could be contrasted with an infinite-far-field boundary condition - which, in accordance with Section 2.2.4, is both possible and likely more performant. However, this explicit freestream boundary formulation is an equally-valid PDE formulation, and it serves as a useful example case here.. Here, the freestream velocity is implemented as a Dirichlet condition on the velocity field at the freestream boundary. Of course, the direction and magnitude of this freestream velocity vector has a massive impact on the near-field flow around the object of interest. Without the hyperlayer architecture, these effects would need to be mediated by the (nonphysical and distant) freestream boundary. Instead, the hyperlayer architecture allows the freestream boundary to propagate its information onto the body boundary, allowing the body boundary in turn to handle near-field effects without high-magnitude, long-range kernel behavior.
3 Modeling Tasks, Datasets, and Training Considerations
We evaluate GLOBE on AirFRANS, a public benchmark by Bonnet et al. (2023) consisting of steady, incompressible RANS solutions over NACA 4- and 5-digit airfoils spanning high-Reynolds, subsonic conditions. The dataset comprises 1,000 simulations with Reynolds numbers in $2,6]\times 10^{6}$ and angles of attack ranging from $5^{\circ}$ to $15^{\circ}$. It is released in both raw OpenFOAM format and a preprocessed VTU/VTP format with clipping to a near-field region, a mid-span slice at $z=0.5$, inward airfoil normals, and a distance function, retaining pressure, velocity, and turbulent viscosity fields; we use the latter. We adopt the four official evaluation regimes defined by the dataset authors:
  • Full (interpolation). 800 simulations for training and 200 for validation/test, drawn i.i.d. from the same distribution.
  • Scarce (low-data interpolation). Same validation/test set as Full but with 200 simulations for training.
  • Reynolds extrapolation. Train on cases with $\rm Re\in[3,5]\times 10^{6}$; evaluate on cases outside this range.
  • Angle-of-attack (AoA) extrapolation. Train on cases with $\alpha\in[-2.5^{\circ},12.5^{\circ}]$; evaluate on the complementary set.
Preprocessing and nondimensionalization.
We use the preprocessed meshes and follow a strictly unitless pipeline implemented in our reference data pipeline. For each sample, we load the internal VTU and the two VTP boundaries (freestream and airfoil) and construct an explicit boundary mesh for the airfoil with a ‘no_slip‘ BC using the face centers and normals. Let $\bm{U}_{\infty}$ denote the mean freestream velocity on the freestream boundary, with magnitude $|\bm{U}_{\infty}|$, density $\rho=1\,\mathrm{kg/m^{3}}$ [^14] While some secondary sources report $\rho=1.204\,\mathrm{kg/m^{3}}$, the OpenFOAM cases in AirFRANS are configured with $\rho=1$. This choice preserves constant far-field total pressure, consistent with incompressible flow assumptions., and kinematic viscosity $\nu=1.56\times 10^{-5}\,\mathrm{m^{2}/s}$. We define the dynamic pressure $q_{\infty}=\tfrac{1}{2}\rho|\bm{U}_{\infty}|^{2}$ and compute the following nondimensional fields on the internal point set:
$\frac{\bm{U}}{|\bm{U}_{\infty}|},\quad\frac{\Delta\bm{U}}{|\bm{U}_{\infty}|}$ $=\frac{\bm{U}-\bm{U}_{\infty}}{|\bm{U}_{\infty}|},$
$C_{p}$ $=\frac{p}{q_{\infty}},\qquad C_{pt}=\frac{p+q}{q_{\infty}},\quad q=q_{\infty}\left\lVert\frac{\bm{U}}{|\bm{U}_{\infty}|}\right\rVert^{2},$
$\ln\bigl(1+\nu_{t}/\nu\bigr)$ $,\qquad\bm{C}_{F,\text{shear}}=\frac{2\nu\,\bm{S}\,\bm{\hat{n}}_{w}}{q_{\infty}},$
where $\bm{S}=\tfrac{1}{2}(\nabla\bm{U}+\nabla\bm{U}^{\!T})$ is the symmetric strain-rate tensor and $\bm{\hat{n}}_{w}$ are per-point airfoil normals sampled at the surface-only points.
Inputs and output fields for GLOBE.
Each sample provides: (i) a list of prediction points (all internal mesh points), (ii) a single boundary mesh representing the airfoil surface with the boundary condition type labeled ‘no_slip’, (iii) global vectors containing $\bm{U}_{\infty}/|\bm{U}_{\infty}|$, and (iv) reference lengths. The model predicts a selected set of nondimensional fields at the prediction points. In our main experiments we predict the vector field $\Delta\bm{U}/|\bm{U}_{\infty}|$ and scalars $C_{p},C_{pt},\ln(1+\nu_{t}/\nu)$. We additionally include a wall-shear vector channel $\bm{C}_{F,\text{shear}}$, which is only interpreted at boundary points.
During training, we mask out all points in the dataset with obviously non-physical total pressures ($C_{pt}>1.02$), as these denote significant violations of energy conservation – this affects 0.003% of points across the entire dataset.
Reference lengths provided to GLOBE are $\ell\in\{c_{\text{ref}},\ \delta_{FS}\}$, where $c_{\text{ref}}$ is the airfoil chord and $\delta_{FS}=\sqrt{\nu\,c_{\text{ref}}/|\bm{U}_{\infty}|}$ is a freestream viscous thickness scale.
On excluding signed distance fields.
While AirFRANS includes signed distance fields (SDFs), we deliberately exclude them as model input features. SDF features pose two fundamental problems in PDE surrogate modeling: (1) SDFs have no direct physical significance in the governing equations - they are purely geometric constructs. If proximity information were physically meaningful, better approaches would use fields derived from physical PDEs, such as the signed heat method of Feng & Crane (2024). (2) More critically, SDFs are inherently non-smooth: gradients exhibit discontinuities at equidistant loci (medial axes/surfaces), violating the $C^{\infty}$ continuity that GLOBE’s kernel architecture maintains (Section 2.2.2). The relative position $\bm{r}_{ts}$ (Eq. 1) provides sufficient geometric information without these pathologies.
On redundant output fields, gradients, and self-consistency.
Several output fields are not independent: for example, given velocity and $C_{p}$, one can infer $C_{pt}$ (and conversely), and wall-shear vectors could be obtained from velocity gradients [^15] Specifically: forward-mode differentiate the velocity field to obtain the velocity gradient tensor, apply the viscosity constitutive relation to compute the stress tensor, then contract with the surface normal vector to extract the wall shear force vector.. Therefore, it would be possible to implement a Sobolev-style objective to enforce these relationships implicitly through the loss.
However, we find that directly predicting both underlying and derived quantities yields measurably better accuracy and faster convergence, mirroring empirical observations in Sobolev training, where such redundancy stabilizes optimization and improves fidelity in sharp-gradient regions (e.g., near-wall shear). A further practical benefit is trustworthiness assessment during extrapolation: redundant channels enable lightweight “self-inconsistency” diagnostics at inference (e.g., comparing the explicitly predicted $C_{pt}$ to the value implied by $\bm{U}$ and $C_{p}$). Elevated self-inconsistency flags untrustworthy predictions, an important consideration for CAE workflows that is often underemphasized in ML surrogate literature. That said, eliminating redundant channels and enforcing physical relationships through exact analytic relations is elegant and could strengthen inductive bias - we leave this as a promising direction for future work.
4 Results and Discussion
For each of the four tasks identified in Section 3, we train a unique GLOBE model, with full training details and hyperparameter values in Appendix E. The following sections discuss the accuracy of these models, both quantitatively and qualitatively.
4.1 Error Metrics
Headline accuracy.
As seen in Tables 3 and 4, GLOBE achieves substantial accuracy improvements over existing models. On the Full split, it reduces mean squared error (MSE) by roughly 200x on all fields relative to the reference baseline model architectures released by the AirFRANS dataset authors, Bonnet et al. (2023). Compared to the next-best-performing model, enf2enf by Catalani et al. (2025a), where only pressure metrics are available in the literature for comparison, we find that GLOBE achieves roughly 50x lower error.
On the Scarce split, where the latest state-of-the-art models have been benchmarked by Catalani et al. (2025b), GLOBE reduces MSE by roughly 7x on velocity and pressure fields and 66x on surface-only pressure when compared to the next-best-performing model, MARIO. Compared to Transolver, a popular model architecture by Wu et al. (2024), GLOBE achieves reductions in MSE of over 100x on velocity and pressure fields and over 600x on surface-only pressure.
Table 3: AirFRANS Full (interpolation) task: validation-set mean squared error on z-score-normalized fields; lower is better. Metrics with (–) were not reported in the associated citation.
Mean Squared Error, “Full” task
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (our work) 0.0047 0.0039 0.0031 0.0039
enf2enf 0.1100 0.3200
CORAL∗∗ 0.3500 1.2000
Transolver 0.3700 1.4200
MeshGraphNet†† 2.1400 3.8700
GINO|| 2.9700 4.8200
MLP 0.9490 0.9780 0.7370 11.3000
GraphSAGE 0.8320 0.9940 0.6610 6.6200
PointNet 3.5000 3.6450 1.1510 9.2500
GUNet 1.5170 2.0280 0.6570 3.8600
Architecture and metrics from Catalani et al. (2025a).
∗∗ Architecture from Serrano et al. (2023), metrics from Catalani et al. (2025a).
Architecture and metrics from Wu et al. (2024).
†† Architecture from Pfaff et al. (2021), metrics from Catalani et al. (2025a).
|| Architecture from Li et al. (2023), metrics from Catalani et al. (2025a).
(AirFRANS dataset reference models) Implementations and metrics from Bonnet et al. (2023).
Table 4: AirFRANS Scarce (low-data) task: validation-set mean squared error on z-score-normalized fields; lower is better.
Mean Squared Error, “Scarce” task
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (our work) 0.0208 0.0180 0.0286 0.0392
MARIO§ 0.1520 0.1130 0.2400 2.7000
Transolver 2.1050 2.1080 4.4340 23.6700
MLP 1.6470 1.4510 3.9040 21.9200
GraphSAGE 1.4570 1.4540 4.6960 19.4500
PointNet 3.1110 2.7760 3.2940 18.2700
GUNet 1.7490 1.8250 3.3880 14.7300
§ Architecture and metrics from Catalani et al. (2025b).
Architecture from Wu et al. (2024), metrics from Catalani et al. (2025b).
(AirFRANS dataset reference models) Implementations and metrics from Bonnet et al. (2023).
Table 5: AirFRANS Reynolds (extrapolation) task: validation-set mean squared error on z-score-normalized fields; lower is better.
Mean Squared Error, “Reynolds” task
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (our work) 0.1897 0.0811 0.2446 0.2903
MLP 9.5050 4.9240 4.3000 208.9800
GraphSAGE 7.5580 3.4980 3.8260 17.9700
PointNet 9.4220 7.1290 4.0110 20.1300
GUNet 8.3830 5.2500 4.4830 20.5900
(AirFRANS dataset reference models) Implementations and metrics from Bonnet et al. (2023).
Table 6: AirFRANS Angle of Attack (extrapolation) task: validation-set mean squared error on z-score-normalized fields; lower is better.
Mean Squared Error, “Angle of Attack” task
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (our work) 0.9706 1.1928 11.3786 21.8215
MLP 6.9650 10.6300 11.7110 87.6200
GraphSAGE 4.4350 9.4000 10.9080 76.3800
PointNet 8.6800 15.7960 16.2370 58.4600
GUNet 5.6890 10.3420 14.8870 69.6700
(AirFRANS dataset reference models) Implementations and metrics from Bonnet et al. (2023).
Nondimensionalization and error metric reporting methodology.
In Tables 36, the notations $\bar{u}_{x}$ and $\bar{u}_{y}$ correspond to z-score-normalized dimensional velocity components, i.e., $(u_{x,\text{pred}}-u_{x,\text{true}})/\sigma$, where $\sigma$ is the standard deviation of $u_{x,\text{true}}$. Likewise, $\bar{p}$ and $\bar{p}_{s}$ refer to z-score-normalized dimensional pressure on the volume and surface, respectively. These metrics exactly match the methodology actually implemented in the AirFRANS reference code [^16] We note a discrepancy regarding normalization methodology in the AirFRANS work: while Bonnet et al. (2023) states that “each quantity is normalized either by $u_{\infty}$ the inlet velocity magnitude or $\nu$ the fluid viscosity,” indicating nondimensionalization based on reference quantities, the released code implementation instead uses z-score statistical normalization. To ensure valid quantitative comparison with prior benchmarks established using this code, we replicate the z-score normalization of the reference implementation here, despite the physical limitations discussed in this section.. While necessary for apples-to-apples comparison with prior work, we emphasize that z-score-based normalization is not recommended for aerodynamic applications, and is at odds with standard practice in fluid dynamics literature.
Z-score-based normalization has three fundamental problems for physics applications: (1) the normalization scale varies across different flow conditions, geometries, and datasets, precluding meaningful cross-study comparisons; (2) it obscures the connection to physical quantities of interest – aerodynamicists care about drag coefficients and pressure recovery in absolute terms, not variance-normalized proxies; and (3) it can artificially inflate apparent performance by down-weighting errors in high-variance regions that may be physically critical (e.g., separated flows, wakes).
The physically principled alternative, and the standard practice in fluid dynamics literature, is nondimensionalization by characteristic scales from the governing equations: velocities by $|\bm{U}_{\infty}|$, pressure by dynamic pressure $\frac{1}{2}\rho|\bm{U}_{\infty}|^{2}$ (yielding the pressure coefficient $C_{p}$), and viscosity by $\nu$ (yielding the turbulent viscosity ratio $\nu_{t}/\nu$). This yields dimensionless quantities with direct physical interpretation that are invariant to dataset statistics.
For completeness, Table 9 in Appendix G reports performance using physical nondimensionalization, which we encourage practitioners to adopt moving forward. These metrics of MSE and MAE on fields such as $\bm{U}/|\bm{U}_{\infty}|$ and $C_{p}$ provide the interpretable, physically-grounded performance assessment that should be standard practice in ML for physics applications.
On reporting $\nu_{t}$.
We do not report error metrics on $\nu_{t}$ because the model is trained on $\ln\left(1+\nu_{t}/\nu\right)$, not on $\nu_{t}$ or the turbulent viscosity ratio $\nu_{t}/\nu$. This is a deliberate choice made for practice-motivated reasons. For incompressible RANS closures, the quantity that actually alters the solution is the effective viscosity $\nu_{\text{eff}}=\nu+\nu_{t}$. Distinguishing $\nu_{t}/\nu$ of 0 vs. 1 (e.g., boundary-layer onset and growth) changes $\nu_{\text{eff}}$ and thus the flow meaningfully, whereas errors between very large values (e.g., 100 vs. 101) in separated recirculation regions have negligible impact on loads or pressure recovery [^17] Recall that the physical purpose behind adding extreme $\nu_{t}$ production in recirculation regions in RANS closures is to capture that recirculation regions are “squishy” (i.e., unable to sustain a pressure gradient, which naturally follows from closed-streamline argument) by causing near-instantaneous diffusion.. Benchmarking error on raw $\nu_{t}$ over-weights these uninformative regions and yields misleading scores. Using $\ln\!\left(\nu_{\text{eff}}/\nu\right)$, or equivalently, $\ln\!\left(1+\nu_{t}/\nu\right)$, linearizes the near-zero regime, compresses the far tail, and aligns the objective with physical relevance; we therefore report metrics on fields that drive forces and pressure recovery.
4.2 Field Visualizations
Figures 47 show GLOBE predictions on validation samples from each of the four tasks, demonstrating sharp boundary-layer gradients, coherent wake structures, and maintained accuracy even under extrapolation. [^18] Validation samples were selected systematically (first in alphabetical order for Full/Scarce; first low-Reynolds and first high-AoA sample for extrapolation splits) rather than cherry-picked for low error. The higher errors in AoA extrapolation compared to Reynolds extrapolation (Table 6 vs. Table 5) likely reflect that angle-of-attack changes induce global shifts in flow topology (e.g., stagnation point movement, wake trajectory), which cause more severe physics changes than the boundary layer thickening associated with Reynolds number changes.
Figure 4: GLOBE model predictions on the first validation sample of the Full split. Rows show ground truth, GLOBE predictions, and error, respectively. Each column corresponds to a different field of interest. For vector fields, the magnitude is shown.
Figure 5: GLOBE model predictions on the first validation sample of the Scarce split. Rows show ground truth, GLOBE predictions, and error, respectively. Each column corresponds to a different field of interest. For vector fields, the magnitude is shown. Even with only 200 training samples, GLOBE results on validation cases qualitatively remain sharp and physically consistent.
Figure 6: GLOBE model predictions on a validation sample from the Reynolds extrapolation split. Rows show ground truth, GLOBE predictions, and error, respectively. Each column corresponds to a different field of interest. For vector fields, the magnitude is shown.
Figure 7: GLOBE model predictions on a validation sample from the AoA extrapolation split. Rows show ground truth, GLOBE predictions, and error, respectively. Each column corresponds to a different field of interest. For vector fields, the magnitude is shown. Even when extrapolating beyond the range of angles of attack seen during training, results remain accurate.
4.3 Training and Inference on Non-Watertight Meshes
A critical challenge in industrial CAE workflows is the preprocessing required to prepare CAD geometry for simulation. In many cases, defeaturing (the manual removal or simplification of small geometric features) is the most time-consuming step in traditional simulation pipelines. Even after defeaturing, practitioners frequently encounter issues with non-watertight meshes: small gaps, overlapping faces, or topological inconsistencies that cause traditional CFD solvers to fail or require further manual intervention. For ML surrogates to achieve practical deployment in industry, they must be robust to such geometric imperfections rather than requiring the idealized meshes typical of academic benchmarks.
The GLOBE architecture is intrinsically well-suited to handling boundary mesh imperfections due to its boundary-element-method-like formulation. Because the kernel evaluation treats each boundary face centroid as an independent source point (Section 2.2), with influences aggregated via area-weighted summation (Eq. 7), the architecture has no explicit dependence on topological connectivity, shared vertices, or edge relationships between faces. Unlike volume-mesh-based methods or graph neural networks that rely on explicit connectivity, GLOBE operates on a cloud of oriented surface elements. This suggests that small gaps, overlaps, or topological defects should degrade performance gracefully rather than catastrophically.
To rigorously test this hypothesis, we conducted an extreme stress test by training on the Full split with severely decimated boundary meshes using a two-step procedure. First, for each training sample, we randomly removed 50% of all airfoil boundary faces. Second, to maintain constant total boundary area (preserving the discretization-invariance property in Eq. 7), we doubled the area of each remaining face by moving its vertices radially outward from the face centroid. Critically, we treat the mesh as an unordered collection of triangular faces with no connectivity information; previously-adjacent faces no longer share edges after expansion.
This procedure introduces two severe pathologies: (1) the random removal creates large gaps, riddling the airfoil surface with holes where faces were deleted; and (2) the geometric expansion of remaining faces causes them to mutually intersect and creates vertex/edge mismatches with neighbors, destroying all topological connectivity. The resulting meshes are geometrically degenerate by any conventional standard; no traditional CFD solver would accept such input without extensive repair.
This test also serves as an extreme empirical validation of the discretization-invariance claim in Section 1.4: despite a radical change in the discrete representation (50% face removal), the integral formulation ensures the solution remains stable.
Despite this severe degradation of input quality, GLOBE produces reasonable predictions. Table 7 compares validation-set performance across three configurations: the standard Full model, the decimated-mesh model, and the Scarce model (included as a data-scarcity reference point). While errors increase under decimation (as expected when discarding half the boundary information), the model remains functional, with performance degradation comparable to the reduction in training data quantity (Scarce regime). Put another way, the error of GLOBE using this heavily-decimated training data is still far lower than the error of all other models from the literature that were compared in Table 3 and Table 4.
Most critically, the model does not collapse or exhibit pathological failures despite being trained exclusively on topologically invalid meshes and evaluated on intact validation meshes, demonstrating substantial robustness to geometric preprocessing artifacts.
Table 7: Robustness to non-watertight meshes: GLOBE trained on severely decimated boundary meshes (50% of faces randomly removed, remaining faces doubled in area via vertex expansion, destroying topological connectivity) compared to baseline Full and Scarce models. Training uses decimated meshes; validation uses intact meshes. Metrics are mean squared error on z-score-normalized fields.
Mean Squared Error
Model $\bar{u}_{x}\;(\times 10^{2})$ $\bar{u}_{y}\;(\times 10^{2})$ $\bar{p}\;(\times 10^{2})$ $\bar{p}_{s}\;(\times 10^{2})$
GLOBE (Full) 0.0047 0.0039 0.0031 0.0039
GLOBE (Full, decimated) 0.0510 0.0402 0.0311 0.0336
GLOBE (Scarce) 0.0208 0.0180 0.0286 0.0392
Figure 8 shows a representative validation sample predicted by the decimated-mesh model. Even when trained exclusively on meshes with 50% missing faces and severe topological defects, the model recovers coherent flow structures, sharp boundary-layer gradients, and physically plausible wakes. This robustness has important practical implications: GLOBE can potentially be deployed on CAD-derived meshes with minor geometric imperfections without requiring the extensive manual cleanup that traditional solvers demand, significantly reducing preprocessing overhead in industrial workflows.
Beyond handling defeaturing artifacts, this decimation robustness also provides a near-term path to scalability for large-scale 3D datasets. The results suggest that training on heavily downsampled boundary meshes does not catastrophically degrade accuracy - a critical property when extending to industrial-scale 3D geometries where boundary meshes can contain millions of faces. While hierarchical acceleration methods (Section 2.2.7) offer a long-term solution to the quadratic complexity of all-to-all kernel evaluation, downsampled training provides an immediately practical approach: we can train on heavily coarsened boundary representations with acceptable accuracy tradeoffs, then evaluate on full-resolution meshes at inference. This contrasts sharply with connectivity-dependent architectures (GNNs, volume-mesh CNNs) where downsampling fundamentally alters the graph structure or requires expensive interpolation, often leading to catastrophic performance collapse. The area-weighted aggregation in Eq. 7 ensures that as long as total boundary area is preserved, the discretization level primarily affects representational fidelity rather than architectural validity.
Figure 8: GLOBE model predictions on a validation sample when trained on severely decimated meshes (50% boundary face removal). Rows show ground truth, GLOBE predictions, and error, respectively. Each column corresponds to a different field of interest. For vector fields, the magnitude is shown. Despite training exclusively on topologically invalid meshes with massive gaps and intersections, the model produces accurate predictions due to its formulation.
5 Computational Reproducibility
All source code used to generate the results in this publication will be available within the open-source NVIDIA PhysicsNeMo framework at https://github.com/NVIDIA/physicsnemo in its subsequent release; in the interim, interested readers may contact the authors to inquire about access.
6 Conclusions
In this work, we show that rigorous incorporation of physical structure and domain-inspired inductive biases can yield substantial improvements in accuracy, generalizability, and practicality for ML-based PDE surrogates. By synthesizing insights from boundary-element methods, exact physical symmetries, and learned kernel representations, GLOBE achieves strong performance on aerodynamic surrogate modeling: roughly $200\times$ error reduction over dataset baselines on interpolation tasks, roughly $10$ to $100\times$ improvement over state-of-the-art methods in low-data regimes, and robust extrapolation under Reynolds number and angle-of-attack distribution shifts. These gains are achieved with a compact architecture (117k parameters) that evaluates fields at arbitrary points, handles non-watertight meshes, and respects the necessary attributes of valid physics models established in Section 1.4.
The core architectural innovations (equivariant kernel functions, Padé-approximant MLPs, multiscale composition, and communication hyperlayers) collectively encode the structural prior that solutions to boundary-driven elliptic PDEs can be represented as superpositions of source influences from boundaries. This boundary-centric formulation proves highly effective when combined with architecturally-enforced translation, rotation, and parity equivariances; discretization-invariance through area-weighted boundary integrals; units-invariance via rigorous nondimensionalization; and global receptive fields that respect elliptic PDE information flow requirements.
Beyond the specific numerical results, this work provides broader lessons about the appropriate role of structure in machine learning for physics. The oft-cited “bitter lesson” of Sutton, which states that general-purpose methods leveraging computation ultimately outperform approaches with hand-coded knowledge, reflects legitimate insights from decades of progress in perception, language, and symbolic reasoning. However, physics presents some important differences. While domains like computer vision and linguistics have structure, exceptions and ambiguous cases are common. In physics, conservation laws, symmetries, and dimensional consistency are exact mathematical constraints that hold universally across problem instances, not approximate patterns to be learned from data.
Empirical evidence supports this distinction. For example, even when using massive datasets like DrivAerML’s 31 terabytes of CFD data (Ashton et al. (2024)), ML surrogate models trained purely from data do not intrinsically learn rotational invariance - a symmetry that follows directly from conservation of angular momentum and is exactly preserved by even the simplest physics solvers. When architectures fail to capture such fundamental properties despite enormous data and compute, it suggests that pure scaling is insufficient; some structural encoding is beneficial or perhaps even necessary.
This is not an argument for abandoning machine learning in favor of traditional physics-only approaches. Physics also contains genuine unknowns where general-purpose function approximators excel when surgically applied - turbulence closure, for instance [^19] This is precisely why GLOBE employs learned Padé-approximant MLPs in its core rather than analytical Green’s functions.. Recent hybrid approaches demonstrate the value of this synthesis: Zhang (2022) achieved substantial gains in integral boundary layer methods by using compact neural networks for specific closures while preserving overall physical structure, and Menter & Matyushenko (2025) showed that learning optimal coefficients for physics-based turbulence models from rich simulation datasets yields substantial accuracy improvements over traditional hand-tuned coefficients from canonical experiments. The most effective path forward appears to involve encoding substantial physical structure for the constraints we know to be exact, while applying machine learning surgically to the components that resist analytical derivation. Achieving this synthesis requires deep domain understanding.
The robustness to severely decimated boundary meshes (Section 4) further illustrates the practical value of physics-inspired design. Traditional CFD solvers and many ML surrogates require carefully preprocessed, watertight geometries - a bottleneck that often consumes more engineering time than the simulation itself. GLOBE’s graceful degradation under geometric imperfections, a consequence of its boundary-element-inspired formulation, suggests a path toward surrogates that can operate directly on CAD-derived meshes with minor defects, dramatically reducing preprocessing overhead in design workflows.
Beyond technical efficacy, the successful integration of ML into engineering workflows depends on a critical sociological factor: trust. The CAE community operates under strict standards of verification and validation (V&V), where adherence to physical laws is a prerequisite for credibility. Practitioners who certify safety-critical systems are skeptical of tools that function as black boxes; a model that violates basic symmetries (e.g., predicting different forces upon coordinate rotation) fails these fundamental validity checks regardless of its aggregate error metrics. By structurally enforcing these exact constraints, physics-informed architectures align more closely with the standards needed for acceptance among practicing CAE engineers. This physical interpretability and behavioral guarantees enable further use of ML-based PDE surrogates as reliable industrial tools.
Looking ahead, we identify several directions for future work with GLOBE:
  • Scaling to 3D aerodynamic datasets (e.g., DrivAerML) to extend the approach beyond 2D airfoil analysis. The decimation results suggest training on downsampled boundary meshes provides a near-term path; hierarchical kernel acceleration (reducing complexity from $\mathcal{O}(n_{\text{src}}^{2})$ to $\mathcal{O}(n_{\text{src}}\log n_{\text{src}})$) would enable longer-term full-resolution training on geometries with hundreds of thousands of boundary faces.
  • Evaluating on other PDE domains beyond aerodynamics (e.g., static structural analysis, heat transfer, electromagnetics) to validate architectural generality.
  • Systematic ablation studies isolating individual contributions of Padé MLPs, far-field decay envelopes, multiscale composition, and communication hyperlayers to strengthen understanding of which innovations drive performance.
  • Mixed-BC problem evaluation (e.g., automotive aerodynamics with no-slip/free-slip boundaries) to demonstrate architectural flexibility in settings requiring explicit BC type encoding, beyond the single-BC problems in the current evaluation.
  • Additional physical structure such as a) divergence- and curl-free conditions on appropriate vector fields, or b) new forms of mass/momentum conservation constraints that avoid the serious pitfalls of PINN-like strategies that use residual-based losses. Whether such constraints improve performance or prove redundant given existing physical structure remains an open question.
We believe that the path to practical surrogates in physics is not only through scaling, but also through hybrid approaches that leverage analytical components where exact structure is known from first principles, and the flexibility of ML methods where it is not. GLOBE demonstrates that substantial gains remain achievable through careful architectural design - and that the pursuit of such physics-informed architectures, far from limiting ML’s flexibility, may be what ultimately enables its full potential in scientific computing.
Acknowledgements
We thank Corey Adams for valuable feedback and paper review, and the NVIDIA CSRG team for maintaining supercomputing resources that enabled this research.
Appendix Appendix A Detailed Background on Integral Boundary Layer Methods
For 2D aerodynamic flows without massive separation, specialized integral boundary layer (IBL) methods such as XFOIL and MSES by Drela (1989; 2007) provide an existence proof that dramatically faster and more accurate methods are possible. IBL methods decompose the flow into a boundary layer and an inviscid outer region, which are then coupled. The key benefit is that the inviscid outer flow can often be modeled as a potential flow, which is a linear, homogeneous PDE. This allows discretization using the enormously efficient boundary-element method (BEM), which discretizes only boundaries rather than the entire volume, yielding 100- to 1,000-fold reductions in both degrees of freedom and computational cost, as shown by Drela (2013b); Martinsson (2020).
Remarkably, IBL methods often achieve higher accuracy than RANS CFD [^20] Even then, for RANS CFD methods to reach accuracy parity with IBL on attached aerodynamic flows with transitional boundary layers, RANS solvers require sophisticated turbulence models like the Re-$\gamma$ SST model by Langtry & Menter (2009) – a critical step that is often overlooked. on their target problems, as shown by Coder & Maughmer (2014); Morgado et al. (2016); Adler et al. (2022). Because of this, IBL methods are the gold standard for attached 2D aerodynamic flows among industry practitioners.
However, extending IBL to 3D remains elusive despite substantial efforts, with notable advances in Nishida (1996); Drela (2013a; b; 2019); Zhang (2017; 2022). Challenges include wake treatment, stability on non-smooth geometries, and the enormous theoretical complexity of 3D IBL equations, as shown in derivations by Drela (2013b).
Nevertheless, IBL’s success in 2D demonstrates that boundary-centric architectural ideas encode powerful inductive biases for fluid dynamics: explicitly decomposing flow into boundary and outer regions, with global information transfer mediated through boundary-only integrals. The GLOBE architecture draws inspiration from these principles while leveraging ML’s flexibility to extend beyond the constraints of analytical IBL methods.
Appendix Appendix B Illustrations of Kernel Functions
In this appendix we visualize an example kernel function $\mathcal{K}$, which aids in demonstrating some of the kernel’s key mathematical properties, as well as its expressive range. While the example shown here is in 2D for ease of visualization, all constructions and guarantees (translation/rotation/parity equivariance, discretization invariance, and far-field behavior) are identical in 3D; see Sections 2.22.3 for descriptions on the minor architectural changes associated with this.
Each figure shows a scalar field $\phi$ and a vector field $\bm{u}$ produced by $\mathcal{K}$ when evaluated from boundary face “sources” to a dense set of “target” points (Eq. 1) in physical space ($\mathbb{R}^{2}$). The axes in all figures represent the $x$- and $y$-coordinates of the target points in nondimensional spatial coordinates (i.e., divided by the reference length $\ell$, as discussed in Section 2.2.1). For simplicity, and to facilitate a later discussion on symmetry, all visualizations use kernel functions that do not include any global scalars or global vectors as inputs; only per-source vectors (when present) are provided. The parameters of the kernel function’s learned Padé-approximant MLP core are their initial pseudorandomly-generated values, and are the same for the kernels across all figures.
Single source at the origin.
We begin with a single source at the origin. The source carries a vector-valued input (e.g., the unit normal of the boundary face, or a vector-valued boundary condition), which participates in the invariant feature construction (Section 2.2.2) and later vector reprojection (Section 2.2.5). Figure 9 shows the resulting fields, with the source denoted by a black dot and its associated vector-valued input denoted by a black arrow. Two asymptotic behaviors are worth noting (Section 2.2.4): (i) self-influence is driven to zero by the $(1-e^{-\left\lVert\bm{r}_{ts}\right\rVert^{2}})$ factor, avoiding overly compact support; and (ii) as $\left\lVert\bm{r}_{ts}\right\rVert\to\infty$, the envelope enforces $\sim 1/r$ decay in 2D and $\sim 1/r^{2}$ in 3D, ensuring bounded, physically plausible far-field influence and stable spatial extrapolation.
Figure 9: Single-source kernel evaluation in 2D. The source (black dot) is located at the origin and carries a vector input (black arrow). The far-field envelope (Section 2.2.4) drives self-influence to zero and enforces inverse-distance decay.
Equivariance under translation and rotation.
If we translate and rotate all inputs (the source location and any associated vectors), the outputs translate and rotate exactly. This follows directly from using only the relative position vector $\bm{r}_{ts}=(\bm{x}_{t}-\bm{x}_{s})/\ell$ (Eq. 1) and invariant scalar features; rotational information is reintroduced solely via local reprojection (Section 2.2.5). Because no global vectors or scalars are present and the single source vector in Figures 910 is vertical, the input set is invariant under the reflection $x\mapsto-x$, and parity equivariance therefore implies mirror-symmetric outputs. Figure 10 shows the same single-source pattern after a rigid transformation of the inputs; the fields are correspondingly transformed, confirming exact translation- and rotation-equivariance.
Figure 10: Equivariance demonstration. Rigidly translating/rotating the inputs (source point and its vector) rigidly translates/rotates the output fields without distortion (Sections 2.2.1 and 2.2.5).
Breaking left-right symmetry with an additional vector.
Up to now, the configuration has one source-associated vector. The fields are left-right symmetric because, with no global vectors or scalars and a vertical source vector, the inputs are invariant under left-right reflection; parity-equivariance then mandates symmetry. To explicitly break this invariance, we provide a second source-associated vector (here oriented toward the right), so the input set is no longer reflection-symmetric. The resulting outputs need not be mirror-symmetric, as shown in Figure 11. This behavior follows from the invariant feature construction together with equivariant vector reprojection; the network receives no absolute coordinate frame, and directionality enters only through bases derived from the input vectors (Section 2.2.5).
Note, however, that the resulting fields of Figure 11 are not simply a superposition of the kernel functions evaluated twice with different source-associated input vectors – the two input vectors mutually interact in a manner that is, in general, nonlinear with respect to the input vectors.
Figure 11: Adding a second source-associated vector breaks left-right symmetry, yielding richer, directional patterns (Section 2.2.5).
Necessary radial symmetry without vector information.
Conversely, if no source-associated vectors are provided, there is no directional information with which to break symmetry – this is a direct consequence of the symmetry principle described by Curie (1894). The only geometric input is $\bm{r}_{ts}$’s magnitude, and the learned core only processes (rotationally-invariant) scalars; therefore both $\phi$ and $\bm{u}$ must be radially symmetric. Figure 12 illustrates the resulting radially-symmetric fields. Once again, this symmetry is a consequence of the model’s invariants rather than an explicit hand-coded constraint.
Figure 12: With no source-associated input vectors, the scalar and vector outputs are necessarily radially symmetric. Directionality cannot be inferred without directional inputs.
Superposition of multiple sources.
The final aggregation over sources (Eq. 7) linearly superimposes influences weighted by per-face strengths $w_{s}$ and face areas $a_{s}$. Figure 13 shows four sources placed along a parametric curve, each with its own associated vector. To preserve total influence, each source’s effective strength is reduced ($\propto 1/4$), mimicking the area scaling one would obtain from a coarser boundary discretization.
Figure 13: Superposition from four sources placed along a parametric curve, with per-source strengths reduced to preserve total influence (cf. area weighting in Eq. 7).
Refinement and discretization invariance.
Refining the parametric sampling to twenty sources (each again with reduced strength to maintain the same total) yields fields that closely match those in Figure 13. This mirrors mesh refinement in practice: as boundary faces are subdivided, areas $a_{s}$ decrease and the weighted sum converges. In the limit of vanishing face areas, the discrete contraction in Eq. 7, $\sum_{s}w_{s}\,a_{s}\,\mathcal{K}_{ts}$, converges to the exact boundary surface integral $\int_{\partial\Omega}w(\bm{x}_{s})\,\mathcal{K}(\bm{x}_{t},\bm{x}_{s})\,\mathrm{d}A_{s}$. Figure 14 shows this convergence in action, empirically illustrating the discretization-invariance discussed in Section 2.2.6.
Figure 14: Refinement of Figure 13 to twenty sources. Despite the increased sampling density, the summed fields converge, demonstrating discretization-invariance under face-area weighting (Section 2.2.6).
Appendix Appendix C Illustrations of Padé-Approximant MLPs
This appendix visualizes the functional behavior of randomly-initialized Padé-approximant MLPs introduced in Section 2.2.3. We use the notation Padé $[N/D]$ to denote a Padé MLP with numerator order $N$ and denominator order $D$, mirroring mathematical convention from traditional polynomial-based Padé approximants.
To illustrate the diversity of functional forms that can be learned by Padé-approximant MLPs, Figure 15 shows 25 random initializations of a Padé $[1/3]$ MLP. For comparison, Figure 16 shows 25 random initializations of a standard MLP with the same layer sizes and activations.
While both types of MLPs satisfy the universal approximation theorem, the classes of functions that these networks most readily express are quite different. We suggest that the functional forms expressed by Padé-approximant MLPs may be better suited for representing typical Green’s functions arising in physics-based PDEs.
Figure 15: 25 random initializations of a $\mathbb{R}^{1}\to\mathbb{R}^{1}$ Padé $[1/3]$ MLP with layer sizes $[1,64,64,1]$ and SiLU activations on each subnetwork, with each network represented as a different line.
Figure 16: 25 random initializations of a $\mathbb{R}^{1}\to\mathbb{R}^{1}$ standard MLP with layer sizes $[1,64,64,1]$ and SiLU activations, with each network represented as a different line.
Appendix Appendix D Formal Proofs of GLOBE’s Mathematical Properties
We now formally establish that GLOBE satisfies the necessary attributes from Section 1.4.
Translation-equivariance.
All positional information enters the kernel through the relative position $\bm{r}_{ts}=(\bm{x}_{t}-\bm{x}_{s})/\ell$ (Eq. 1). Under translation $\bm{x}\mapsto\bm{x}+\bm{c}$, we have $\bm{r}_{ts}\mapsto(\bm{x}_{t}+\bm{c})-(\bm{x}_{s}+\bm{c})/\ell=\bm{r}_{ts}$, unchanged. All subsequent operations depend only on $\bm{r}_{ts}$ and other translation-invariant quantities (normals, boundary data), so translating all inputs translates outputs identically.
Rotation-equivariance.
The kernel processes only rotationally-invariant scalars: magnitudes $\operatorname{smoothlog}(\left\lVert\bm{v}_{i}\right\rVert)$ and pairwise angles via Legendre polynomials. Physical vectors are reconstructed by reprojecting onto local bases derived from input vectors (Section 2.2.5). Under rotation $\mathcal{R}$, each input vector transforms as $\bm{v}_{i}\mapsto\mathcal{R}\bm{v}_{i}$, but magnitudes and angles remain invariant. The basis vectors transform as $\bm{b}_{j}\mapsto\mathcal{R}\bm{b}_{j}$, and the reprojection $\sum_{j}c_{j}\bm{b}_{j}\mapsto\sum_{j}c_{j}(\mathcal{R}\bm{b}_{j})=\mathcal{R}\left(\sum_{j}c_{j}\bm{b}_{j}\right)$ yields exact equivariance. The final calibration (linear for vectors) preserves this property.
Parity-equivariance.
The architecture respects parity when tangential/vortex basis vectors $\bm{e}_{\theta}$, $\bm{e}_{\phi}$ are disabled (these are pseudovectors), as is the default. With only true vectors in the reprojection basis (radial, axis, dipole directions), all bases transform correctly under reflection, and the Legendre polynomial features are parity-even. Full parity preservation including tangential bases requires tracking vector vs. pseudovector semantics (Section 2.2.5); this is left as future work.
Discretization-invariance.
The aggregation in Eq. 7 weights each source by its face area $a_{s}$. As the boundary mesh is refined with $a_{s}\to 0$, the discrete sum $\sum_{s}w_{s}a_{s}\mathcal{K}_{ts}$ converges to the continuous boundary integral $\int_{\partial\Omega}w(\bm{x}_{s})\mathcal{K}(\bm{x}_{t},\bm{x}_{s})\,\mathrm{d}A_{s}$. The kernel itself is mesh-independent (parameterized by continuous positions), so predictions converge to a well-defined limit as discretization is refined.
Units-invariance.
All positional features are nondimensionalized by reference lengths $\ell_{i}$. All input scalars and vectors must be dimensionless, a requirement that should be enforced during data preprocessing. Our implementation for the case study in Section 3 demonstrates one approach to ensuring this property. Kernel operations preserve dimensionlessness throughout, and physical units are recovered only during post-processing. Therefore, the entire forward pass operates in a unitless regime, guaranteeing units-invariance.
Appendix Appendix E Hyperparameters and Training Details
Implementation note: TensorDict data structure.
[…]
Training protocol.
During training, we employ dynamic subsampling of prediction points: on each iteration, a limited number of prediction points are randomly chosen to be the query points for training (default: 4,096). The loss is a modified [^22] Naïve use of a Huber objective on individual components of a vector field breaks rotational invariance of the loss function; we correct for this. per-field Huber objective [^23] Chosen to minimize the influence of any remaining nonphysical artifacts in the training data. with scales $\left\{\Delta U/|U_{\infty}|:1,\ C_{p}:1,\ C_{pt}:1,\ \ln(1+\nu_{t}/\nu):5,\ C_{F,\text{shear}}:10^{-2}\right\}$, applied after masking invalid ground-truth entries. Optimization uses a combined optimizer: Muon by Jordan et al. (2024) for matrix-shaped parameters and RAdamW by Liu et al. (2021) for all other parameters, with ReduceLROnPlateau scheduling. We train in bfloat16 with automatic mixed precision [^24] In the bfloat16 case, note that the aggregation step within kernel functions benefits from localized up-casting to float32 to mitigate catastrophic cancellation. and use torch.compile in the "max-autotune" mode (i.e., including CUDA Graphs). Distributed data-parallel training is used when multiple GPUs are available, and parallelization experiments show near-linear scaling of train-time effective throughput with increasing accelerator count, even at the largest-benchmarked configuration with 128 GB200 GPUs. Validation follows the official split per regime and uses the official prediction set (i.e., without any subsampling or masking) for both the accuracy metrics and qualitative visualizations in Section 4.
Table 8 lists all hyperparameters used in the main AirFRANS experiments. These values were selected through systematic but limited exploration: we compared 2-layer (128 neurons) versus 3-layer (64 neurons) architectures, swept learning rate over $10^{-2}$, $10^{-3}$, $10^{-4}$}, and explored three weight decay values. No extensive hyperparameter tuning was performed; the reported configuration represents a reasonable baseline rather than an exhaustively optimized setup, suggesting that further improvements may be possible with more thorough search. In total, these hyperparameters lead to a model with 117,282 trainable parameters.
Table 8: Hyperparameters for AirFRANS experiments.
Hyperparameter Value
Architecture
   Number of communication hyperlayers ($H$) 2
   Hidden layer sizes [64, 64, 64]
   Number of latent scalars per hyperlayer 6
   Number of latent vectors per hyperlayer 3
   Number of spherical harmonics ($n_{\text{harmonics}}$) 1
   Padé numerator order ($N$) 2
   Padé denominator order ($D$) 2
   Reference lengths $\{c_{\text{ref}},\delta_{FS}\}$
Training
   Query points per iteration 4,096
   Boundary face downsampling ratio 1.0 (No downsampling)
   Learning rate (base) $10^{-3}$
   Weight decay $10^{-4}$
   Scheduler ReduceLROnPlateau
    Patience 400 epochs
    Factor 0.5
    Min LR $6.25\times 10^{-5}$
   Optimizer Muon (matrices) + RAdamW (others)
   Precision bfloat16 (with AMP)
   Huber loss delta 1.0
Appendix Appendix F Computational Efficiency
Beyond accuracy, computational efficiency is critical for practical deployment. We analyze the scaling characteristics of GLOBE during both training and inference.
Training efficiency.
Training was performed using distributed data parallelism, with linear throughput scaling observed when tested up to the largest-benchmarked configuration of 128 NVIDIA GB200 GPUs. This efficient scaling is enabled by the dense compute intensity of the $\mathcal{O}(n_{\text{src}}^{2})$ boundary-to-boundary interactions, which dominates communication overhead.
With a training configuration of 40 GB200 GPUs, using the Full split dataset described in Section 3 and the hyperparameters listed in Table 8, we record an epoch loop time of 6.83 seconds. With 800 training samples and 40 GPUs, this corresponds to a training speed of 0.341 GPU-seconds per training sample.
Validation samples evaluated for metrics computation are substantially faster, as this consists solely of a forward pass with no need to save primals. Using the same methodology as above, we record a validation speed of 0.083 GPU-seconds per validation sample.
Inference efficiency.
Inference performance is substantially lower than validation performance, as the number of query points is equal to the number of vertices in the volume mesh (approximately $180\times 10^{3}$ points), rather than allowing downsampling. Because of this, we record an inference speed of 2.326 seconds per inference sample on a single GB200 GPU.
Inference memory footprint can be driven down to very low levels by chunking over query points. With a chunk size of 128 points, peak memory consumption is held to approximately 700 MB, allowing deployment on consumer-grade hardware.
Appendix Appendix G Physically Nondimensionalized Performance Metrics
As discussed in Section 4.1, the primary results in Tables 34 use z-score-based normalization to enable direct comparison with prior work on AirFRANS. However, this approach obscures physical meaning. Here we report performance using proper physical nondimensionalization, where velocity fields are normalized by the freestream velocity magnitude $|\bm{U}_{\infty}|$, pressure by the dynamic pressure $\frac{1}{2}\rho|\bm{U}_{\infty}|^{2}$ (yielding the pressure coefficient $C_{p}$), and turbulent viscosity by the kinematic viscosity $\nu$.
Table 9 presents mean squared error (MSE) and mean absolute error (MAE) on both the Full and Scarce validation sets using these physically meaningful nondimensional quantities. Unlike Tables 34, values are reported without the $\times 10^{2}$ multiplier (which was used in the main tables only for consistency with AirFRANS reporting style). We report MAE as the primary metric: it provides direct physical interpretability in the base units of each field. For example, an MAE of $2.1\times 10^{-3}$ on $C_{p}$ indicates that typical errors are $\pm 0.0021$ in pressure coefficient – immediately meaningful to practitioners. In contrast, MSE values require taking the square root for physical interpretation, and overweight outliers in ways that may not align with engineering significance.
Table 9: GLOBE model performance on the AirFRANS validation set, reported using physical nondimensionalization. All fields are nondimensionalized by physical characteristic scales ($|\bm{U}_{\infty}|$ for velocity, $\frac{1}{2}\rho|\bm{U}_{\infty}|^{2}$ for pressure). These metrics provide physically interpretable performance assessment, in contrast to the z-score-based metrics in Tables 34.
Full Dataset Scarce Dataset
Field MSE MAE MSE MAE
Nondimensional velocity
$U_{x}/|\bm{U}_{\infty}|$ $9.53\times 10^{-6}$ $1.57\times 10^{-3}$ $4.53\times 10^{-5}$ $2.38\times 10^{-3}$
$U_{y}/|\bm{U}_{\infty}|$ $5.45\times 10^{-6}$ $1.14\times 10^{-3}$ $4.62\times 10^{-5}$ $2.11\times 10^{-3}$
Pressure coefficient
$C_{p}$ $3.49\times 10^{-5}$ $2.07\times 10^{-3}$ $1.30\times 10^{-3}$ $5.02\times 10^{-3}$
$C_{p}$ (surface only) $1.06\times 10^{-4}$ $3.71\times 10^{-3}$ $6.04\times 10^{-3}$ $1.09\times 10^{-2}$
$C_{pt}$ $2.59\times 10^{-5}$ $1.73\times 10^{-3}$ $5.19\times 10^{-4}$ $3.41\times 10^{-3}$
$C_{pt}$ (surface only) $1.54\times 10^{-4}$ $5.08\times 10^{-3}$ $8.32\times 10^{-3}$ $1.47\times 10^{-2}$
Turbulent viscosity ratio
$\ln(1+\nu_{t}/\nu)$ $1.97\times 10^{-3}$ $8.30\times 10^{-3}$ $6.24\times 10^{-3}$ $1.40\times 10^{-2}$
Skin friction coefficient (surface only)
$C_{F,\text{shear},x}$ $6.88\times 10^{-9}$ $3.38\times 10^{-5}$ $5.82\times 10^{-8}$ $6.92\times 10^{-5}$
$C_{F,\text{shear},y}$ $1.38\times 10^{-8}$ $3.94\times 10^{-5}$ $1.40\times 10^{-7}$ $9.10\times 10^{-5}$
Interpretation.
[…]
In contrast, the z-score-based metrics in Tables 34 obscure this physical interpretation: without knowing dataset statistics, one cannot infer whether a given error magnitude corresponds to 0.1%, 1%, or 10% physical error. This demonstrates why physical nondimensionalization should be standard practice for ML surrogates in engineering domains, and is indeed longstanding practice in the fluid dynamics literature.